OpenStudy (anonymous):

Susan' rope is three times as long as Fred's rope. Fred's rope is twice as long as Alison's rope. The total length of all three ropes is 36 meters. How long is Fred's rope?

OpenStudy (bahrom7893):

S = 2F; F = 2A S + F + A =36

OpenStudy (bahrom7893):

F = 2A and S = 2F => S = 4A

OpenStudy (anonymous):

Don't understand

OpenStudy (bahrom7893):

S = 4A, F = 2A 4A + 2A + A = 36

OpenStudy (bahrom7893):

Okay so let S be the length of sally's rope

OpenStudy (bahrom7893):

F - fred's A - alison's

OpenStudy (anonymous):

okay

OpenStudy (bahrom7893):

Woops meant not S = 2F, its S = 3F (Three times as long as fred's)

OpenStudy (bahrom7893):

S = 3F F = 2A A S + F + A = 36 (total length)

OpenStudy (bahrom7893):

S = 3F, but F = 2A => S = 3*2A = 6A

OpenStudy (bahrom7893):

So you have: S = 6A F = 2A A = A (A is just A) S + F + A = 36 6A + 2A + A = 36

OpenStudy (bahrom7893):

9A = 36 = > A = 4

OpenStudy (bahrom7893):

F = 2A = 2*4 = 8 meters => Fred's rope is 8 meters long, please click on become a fan if I helped, thanks!

OpenStudy (anonymous):

WHAT IS SUSAN'S LENGTH thANK yOU sUPERHERO, WILL DO

OpenStudy (bahrom7893):

Susan's: S = 6A = 6 * 4 = 24 meters

OpenStudy (anonymous):

hOW DO i BECOME A fAN

OpenStudy (bahrom7893):

Click on Become a fan next to my name (in blue)